OpenStudy (anonymous):

4y-16+8y= -4 What does Y equal?

OpenStudy (anonymous):

add 16 to both sides 4y+8y=16-4=12 add like terms in the left hand side 12y=12 can u do it from here?

OpenStudy (anonymous):

let me try

OpenStudy (anonymous):

WHAT?

OpenStudy (anonymous):

12y=-4+16 12y=12 y=1

OpenStudy (anonymous):

Thankyou ktnguyen1

OpenStudy (anonymous):

you are welcome